OpenStudy (anonymous):

Write an equation in standard form of the line that passes through the given points. (0, 3), (2, –3)

OpenStudy (cathyangs):

standard form is Ax+By=C You can take the slope, y1-y2/x1-x2, and apply it and one point to the point-slope form: (Y-y1)=m(X-x1) with some algebra, you can get this to standard. and in standard, the A has to be a positive number.

OpenStudy (anonymous):

Thanks . :)

OpenStudy (cathyangs):

Glad to help! :D

OpenStudy (anonymous):

so it would be m = -3-3/2-0 = -6/2 ?